OpenStudy (anonymous):

Write y=2x^2+12x+14 in vertex form

OpenStudy (boldjon):

the vertex is at

OpenStudy (anonymous):

sorry am confused how did u get 1/9

OpenStudy (loser66):

factor 2 out, you get y = 2(x^2 +6x +7) inside the parenthesis only: x^2 +6x +7 = x^2 +2*x*3 +3^2 -3^2 +7 = (x+3)^2 -2 Now, put 2 in the original in \(y = 2((x+3)^2 -2) = 2(x+3)^2 -4 \) Make a comparison with the general form: \(y = \color{red}{a}(x- \color{blue}{h})^2 +\color{green}{k}\\y = \color{red}{2}(x- \color{blue}{(-3)})^2 +\color{green}{(-4)}\)